Compare the graphs of y = 2x2 + 8x + 8 and y=x2 without graphing the functions. How can factoring help in comparing the parabolas? Explain in detail.






1
Expert's answer
2021-01-26T17:59:21-0500
"y=2x^2+8x+8=2(x^2+4x+4)=2(x+2)^2"

"f(x)=x^2"

Shift the graph to the left 2 units


"f(x+2)=(x+2)^2"

Stretch vertically by 2


"2f(x+2)=2(x+2)^2"

We can ontain the graph of "y=2x^2+8x+8" from the graph of "y=x^2" using the Graph Transformations.